OpenStudy (anonymous):

Find the slope-intercept equation of the line with the following properties parallel to the 5x+2y=1; containing the point (-5,4).

OpenStudy (anonymous):

to find the slope, lets isolate y 5x + 2y = 1 2y = - 5x + 1 y = -5/2x + 1/2 here the slope is -5/2, and if you are looking for a parallel line, the slope will be the same. So the slope is - 5/2 Now we will put our info in y = mx + b using slope -5/2 and points (-5,4) y = mx + b 4 = -5/2(-5) + b 4 = 25/2 + b 4 - 25/2 = b 8/2 - 25/2 = b - 17/2 = b so your equation is y = - 5/2x - 17/2